OpenStudy (anonymous):

Where does the graph of y = sin x from x = 0 to x = 2π start? somebody please help :(

OpenStudy (bohotness):

what do you think it is ?

OpenStudy (anonymous):

I have no idea, that why I need help

OpenStudy (bohotness):

okay i help you

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

the option are a. at its minimum b.at an x-intercept c. at its maximum d. between an x-intercept and its maximum but I don't know how to get to the answer, and that's what I need help with

jimthompson5910 (jim_thompson5910):

when you replace x with 0, in the equation y = sin(x), what do you get?

jimthompson5910 (jim_thompson5910):

specifically, what is the value of y when x = 0 ?

OpenStudy (anonymous):

uh, I have no idea

jimthompson5910 (jim_thompson5910):

y = sin(x) y = sin(0) ... replace x with 0 y = ??? ... use a calculator to evaluate the sine of 0

OpenStudy (anonymous):

its 0

jimthompson5910 (jim_thompson5910):

correct, y = 0 when x = 0

jimthompson5910 (jim_thompson5910):

so because y = 0, we know that there is an x-intercept here x-intercepts occur when y = 0

OpenStudy (anonymous):

so it meets at its x-intercept? Or inbetweeen the x and its maximum?

jimthompson5910 (jim_thompson5910):

it meets at its x-intercept since it crosses the x axis (y = 0 when that happens) all x-intercepts are points on the curve and on the x axis

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):